# Exceptional Tannaka groups only arise from cubic threefolds

### v2

@@ 6.3. The Fano variety of singular cubic hypersurfaces
By lemma 6.11 it remains to treat the case when $X$ is not normal. We then conclude by proposition 5.3 and examples 6.6, LABEL:, 6.4, LABEL: and 6.8. ∎
6.4. Positivity of Fano surfaces of cubic threefolds
Let $Y\subset\mathbb{P}^{4}$ be a smooth cubic threefold, and $X$ the Fano surface of lines on $Y$. We fix an Albanese embedding $X\hookrightarrow A:=\operatorname{Alb}(X)$. The subvariety $X\subset A$ is nondegenerate, but in general its normal bundle $\mathcal{N}$ fails to be ample. Since $\mathcal{N}$ is globally generated, the ampleness of $\mathcal{N}$ is equivalent to the finiteness of the Gauss map
$\gamma\colon\quad\mathbb{P}(\mathcal{C})\;\longrightarrow\;\mathbb{P}((\operatorname{Lie}A)^{\vee})$
where $\mathcal{C}=\mathcal{N}^{\vee}$ is the conormal bundle [Laz04, ex. 6.1.5]. On the other hand, the hypersurface $Y$ can be recovered from $X$ as the image of the morphism
$\pi\colon\quad\mathbb{P}(T_{X})\;\longrightarrow\;\mathbb{P}(\operatorname{Lie}A)$
for suitable coordinates on $\operatorname{Lie}A$. Via this identification, any line in $Y$ is the image of $\mathbb{P}(T_{X,x})$ for a unique $x\in X$. For any point $p\in\mathbb{P}((\operatorname{Lie}A)^{\vee})$, let $H_{p}\subset\mathbb{P}(\operatorname{Lie}A)$ denote the corresponding hyperplane. Then by the above discussion, the following are equivalent:
  • the fiber $\gamma^{-1}(p)$ is finite, and
  • the cubic surface $Y\cap H_{p}$ contains only finitely many lines.
Note that in any case, we have:
Lemma 6.12.
For any hyperplane $H\subset\mathbb{P}^{4}$ the surface $H\cap Y$ is integral and normal.
Proof.
If the cubic surface $S:=H\cap Y$ is not integral, it contains a plane $P$. Since the dual plane $P^{\vee}$ parametrizes lines on $P$, we have a closed embedding $P^{\vee}\hookrightarrow X$ which is necessarily an isomorphism because $X$ is a surface. This is absurd since for instance $X$ is of general type. Similarly, if $S$ were nonnormal, then the Fano variety of lines $F_{S}$ on $S$ would be of dimension $3$ by examples 6.4, LABEL: and 6.6. But $F_{S}$ embeds into $F_{Y}=X$ which is of dimension $2$, a contradiction. ∎
Lemma 6.13.
For a point $p\in\mathbb{P}((\operatorname{Lie}A)^{\vee})$ the following are equivalent:
  1. (1)
    the fiber $\gamma^{-1}(p)$ is not finite;
  2. (2)
    the cubic surface $H_{p}\cap Y\subset H$ is a cone over an elliptic curve in $\mathbb{P}^{2}$.
Proof.
For (1) $\Rightarrow$ (2) note that the cubic surface $H_{p}\cap Y$ is normal by lemma 6.12, and the only normal cubic surfaces containing infinitely many lines are cones over elliptic curves [BW79, Sak10]. The converse (2) $\Rightarrow$ (1) is clear by example 6.4. ∎
As an immediate consequence, we obtain the following characterization:
Proposition 6.14.
The following are equivalent:
  1. (1)
    The normal bundle of $X\subset\operatorname{Alb}(X)$ is not ample.
  2. (2)
    There is a hyperplane $H\subset\mathbb{P}^{4}$ such that the cubic surface $H\cap Y\subset H$ is a cone over an elliptic curve in $\mathbb{P}^{2}$.
The condition (2) in the above statement occurs if and only if for a suitable choice of homogeneous coordinates on $\mathbb{P}^{4}$ the equation of $Y\subset\mathbb{P}^{4}$ takes the form
[…]
where $f$ is a homogeneous cubic that defines an elliptic curve $E\subset\mathbb{P}^{2}$ and $q$ is a quadratic form. We also remark that the cubic threefolds in question are parametrized by a $32$-dimensional subvariety of $\mathbb{P}(\operatorname{Sym}^{3}(k^{5})^{\vee})\simeq\mathbb{P}^{34}$. In particular, if $Y$ is a generic cubic threefold, then $\mathcal{N}$ is ample. In general, we have the following:
Corollary 6.15.
The morphism $\gamma$ is finite outside a finite set in $\mathbb{P}((\operatorname{Lie}A)^{\vee})$.
Proof.
If the normal bundle $\mathcal{N}$ is ample, the morphism $\gamma$ is finite. Henceforth suppose that $\mathcal{N}$ is not ample and let $p\in\mathbb{P}((\operatorname{Lie}A)^{\vee})$ be a point such that $H_{p}\cap Y$ is a cone over an elliptic curve $E_{p}$. Note that the cubic surface $H_{p}\cap Y$ determines the hyperplane $H_{p}$ uniquely as its linear span. Now each point of $E_{p}$ gives a line on the threefold, hence we have a closed embedding $E_{p}\hookrightarrow X$. Since a nondegenerate surface $X\subset A$ cannot be covered by translates of elliptic curves, there are at most finitely many possible $E_{p}\hookrightarrow X$. But $E_{p}$ determines the cubic surface $H_{p}\cap Y$ by taking the union of all the lines parametrized by $E_{p}$. Since the cubic surface determines $H_{p}$, this leaves only finitely many possibilities for $p$. ∎
Remark 6.16.
The previous statement corrects an error in the remark after (iii) on p. 1462 in [Krä20] where it had been assumed that the Gauss map for the Fano surface is a finite morphism. However, the result in loc. cit. is still true with the same proof, since it only uses [Krä21, lemma 3.3.1] which can still be applied when $\gamma$ is finite over the complement of finitely many points.
@@ 6.5. A construction
Let $X$ be a smooth projective variety over $k$$V$ a finite-dimensional $k$-vector space and
@@
$\textup{T}_{x}\gamma\colon\quad\textup{T}_{x}X\;\longrightarrow\longrightarrow\;\textup{T}_{\gamma(x)}F.$
@@
As a sanity check, note that for $X=\mathbb{P}^{1}$ the hypotheses of proposition 6.17 are never fulfilled, hence this is not in contradiction to proposition 6.1.
@@ 7.1. Main result
    and the difference morphism $d\colon X\times X\to X-X$ has generic degree $\geqslant 6$.
@@
Combining this result with corollary 4.7, we obtain the following classification of subvarieties with Tannaka group $E_{6}$:
@@
Let $X\subset A$ be a smooth irreducible subvariety with ample normal bundle and dimension $<g/2$. Then the following are equivalent:
@@
The rest of the section is devoted to the proof of theorem 7.1. By the Lefschetz principle and [JKLM23, cor. 4.4] we may assume $k=\mathbb{C}$. In this case we already know from section 4.6 that (1) $\Rightarrow$ (2), and (3) $\Rightarrow$ (1) is clear from the example in Example of the introduction. So the only remaining point is (2) $\Rightarrow$ (3), which gives a numerical characterization of Fano surfaces of smooth cubic threefolds.
@@ 7.2. Strategy of the proof
Before coming to the details, let us outline the main steps in the proof of (2) $\Rightarrow$ (3). The idea is to recover the cubic threefold as in the example in Example. By assumption, the subvariety $X\subset A$ is a surface with
@@
Borrow notation from section 6.5 with the short exact sequence (6.1) being (7.1) and $V=\operatorname{Lie}A$. Since $X$ is nondivisible, it is not invariant by translation under any nonzero abelian variety, thus the Gauss map $\gamma\colon X\to\operatorname{Gr}_{2}(\operatorname{Lie}A)$ is a finite morphism [Deb95, prop. 3.1]. By interpreting $c_{2}(\mathcal{N})$ as the top Segre class of $T_{X}$, the positivity of $c_{2}(\mathcal{N})$ implies that the morphism $\pi\colon\mathbb{P}(T_{X})\to\mathbb{P}(\operatorname{Lie}A)$ is generically finite onto its image. The bootstrap result is the computation of the degree of $\pi$. To do this, we use that the difference morphism
@@
As in section 6.5 let $Y\subset\mathbb{P}(\operatorname{Lie}A)$ be the image of $\pi$. Then $Y$ has dimension $3$ and degree
@@
We are left with showing that $Y$ is smooth. We do this by comparing the Hilbert polynomials (with respect to the Plücker embedding) of $X$ and of the Fano variety $F$ of lines on $Y$, which will show:
@@
The rest of the section is devoted to the proof of lemmas 7.3, LABEL:, 7.4, LABEL: and 7.5.
7.3. The projection to the tangent space: Proof of lemma 7.3
Since there is no risk of confusion here, let us write $d=d_{Z}$ and $\pi=\pi_{Z}$. Let $D:=Z-Z$ be the image of the difference morphism. Consider the blow-up $\tilde{A}$ of $A$ in 00, the blow-up $\tilde{D}$ of $D$ in 00 and the blow-up $\tilde{P}$ of $P=Z\times Z$ in the diagonal. The normal bundle of the diagonal in $Z\times Z$ is by definition the tangent bundle of $Z$. It follows that the exceptional divisor of $\tilde{P}\to P$ is $\mathbb{P}(T_{Z})$. On the other hand the exceptional divisor of $\tilde{A}\to A$ is $\mathbb{P}(\operatorname{Lie}A)$ and that of $\tilde{D}\to D$ is the tangent cone $C$ of $D$ at 00. Let $\tilde{d}\colon\tilde{P}\to\tilde{D}$ be the morphism induced by the difference map. The situation is resumed in the following commutative diagram:
@@
To conclude the proof of lemma 7.3, notice first that the top-left square in diagram (7.3) is cartesian. Indeed, this boils down to the fact that the scheme-theoretic fiber of $d$ in 00 is the diagonal $\Delta_{Z}\colon Z\to Z\times Z$. To see this, notice that the difference morphism $A\times A\to A$ is smooth and the scheme-theoretic preimage of 00 is the diagonal $\Delta_{A}\colon A\hookrightarrow A\times A$ with its reduced structure; the claim then follows because the following square is cartesian:
@@
The locus of $T$ where $f$ is finite is open and meets $T^{\prime}$. By replacing $T^{\prime}$ with such an open subset, we may assume that $f$ is finite. Now, degree of $f$ is computed as the generic rank of the coherent $\mathcal{O}_{T}$-module $\mathcal{F}=f_{\ast}\mathcal{O}_{S}$. Let $f^{\prime}\colon S^{\prime}\to T^{\prime}$ be the morphism induced by $f$. Since the morphism $f$ is affine, we have
$\mathcal{F}_{\rvert T^{\prime}}=f^{\prime}_{\ast}\mathcal{O}_{S^{\prime}}.$
The subvariety $S^{\prime}\subset S$ is assumed to be integral, thus its scheme-theoretic image is $T^{\prime}_{{\mathrm{red}}}\subset T^{\prime}$. It follows that the coherent $\mathcal{O}_{T^{\prime}}$-module $f^{\prime}_{\ast}\mathcal{O}_{S^{\prime}}$ is supported in $T^{\prime}_{{\mathrm{red}}}$. Therefore, there is no loss of generality in assuming that $T^{\prime}$ is reduced, hence integral. By the above discussion, the generic degree of $f^{\prime}$ is the generic rank of $\mathcal{F}_{\rvert T^{\prime}}$. The lower semicontinuity of the function $T\ni t\mapsto\operatorname{rk}\mathcal{F}_{t}$ gives the result. ∎
[…]
7.4. The cubic hypersurface: Proof of proposition 7.4
@@
We start by excluding the case $\deg Y=2$. In this case we have $\operatorname{codim}Y=1$ hence $g=5$ and $Y$ is a $4$-dimensional quadric. To rule out the case, let $F$ be the Fano variety of lines in $Y$ and recall by proposition 6.17 that the image of $\gamma\colon X\to F$ is a $2$-dimensional irreducible component of $F$. The rank $r$ of the quadric $Y$ is necessarily 33334444 or $5$ because $Y$ is integral. We proceed by distinguishing case by case:
  • If $r=3$, then $Y$ is a cone over a smooth plane conic $Q$ and $F$ is irreducible of dimension $\dim Q+2(5-3-1)=3$ by proposition 6.3. Contradiction.
  • If $r=4$, then $Y$ is a cone over the normal rational scroll $S(1,1)=\mathbb{P}^{1}\times\mathbb{P}^{1}$ embedded in $\mathbb{P}^{3}$ via the Segre embedding [Har92, p. 285]. By proposition 6.1 the Fano variety of lines in $S(1,1)$ is $\mathbb{P}^{1}\sqcup\mathbb{P}^{1}$, thus $F$ has two irreducible components of dimension $\dim\mathbb{P}^{1}+2(5-4)=3$ by proposition 6.3. This is a contradiction.
@@
The inequality $\deg Y\geqslant 1+\operatorname{codim}Y$ implies that $Y\subset\mathbb{P}(\operatorname{Lie}A)$ has codimension $1$ or $2$. It remains to rule out the codimension $2$ case. If this is the case, the variety $Y$ is of minimal degree, hence by [EH87, th. 1] it is either a cone over a smooth rational scroll or a cone over the Veronese surface, that is, $\mathbb{P}^{2}$ embedded in $\mathbb{P}^{5}$ by the second Veronese embedding. Now taking a cone with respect to some linear projection preserves the codimension. This rules out at once the case of the Veronese surface since it has codimension $3$ while $Y$ has codimension $2$. Suppose $Y$ is a smooth rational normal scroll $S=S(a)$ or a cone over such a scroll. By proposition 6.17 the scheme-theoretic image of the finite morphism $\gamma\colon X\to F$ is one of the irreducible components of $F$. A case by case distinction shows that this is a contradiction. More precisely, the tuple $a=(a_{1},\dots,a_{r})$ satisfies $1\leqslant a_{1}\leqslant\cdots\leqslant a_{r}$ and $a_{1}+\cdots+a_{r}=3$. This leaves us with the $3$ possibilities described in detail in example 6.2:
@@
Suppose $\dim S=1$. In this case the scroll $S$ is the image of the triple Veronese embedding $\mathbb{P}^{1}\hookrightarrow\mathbb{P}^{3}$. As $S$ contains no lines, proposition 6.3 implies that $F$ is irreducible of dimension $3$. But $X$ dominates a component of $F$. Contradiction.
@@
Suppose $\dim S=2$. In this case the Fano variety is $\mathbb{P}^{1}\sqcup\{\textup{pt}\}$. By proposition 6.3 the component of $F$ over $\mathbb{P}^{1}$ has dimension $\dim\mathbb{P}^{1}+2(6-5)=3$ hence $X$ cannot dominate it. Instead again by proposition 6.3 the component of $F$ over the singleton is a projective plane embedded linearly via the Plücker embedding. If $X$ were to dominate it, we would have $\deg\gamma=45$, contradicting (7.2).
@@
Suppose $\dim S=3$. In this case $Y=S$ and $F\simeq(\mathbb{P}^{1}\times\mathbb{P}^{2})\sqcup\mathbb{P}^{2}$. It follows that $X$ has to dominate the component isomorphic to the projective plane. But this component is linearly embedded via the Plücker embedding. As above, if $X$ were to dominate this component, then $\deg\gamma=45$ which contradicts (7.2). ∎
7.5. The Fano surface: Proof of proposition 7.5
@@
According to proposition 6.9 we have $\dim F\leqslant 3$ with equality if and only if $Y$ is either the twisted plane in example 6.8, or a cone over a nonnormal cubic surface or a cone over a smooth cubic plane curve. These cases are excluded by looking at the Fano variety:
@@
  • If $Y$ is a cone over an integral cubic curve, then $F$ is irreducible and has dimension $3$; see example 6.4.
@@
  • If $Y$ is a cone over a nonnormal cubic surface that is not itself a cone over a nonnormal cubic curve, then the $2$-dimensional irreducible components of $F$ are linearly embedded via the Plücker embedding; see example 6.6. This would imply $\deg\gamma=45$ contradicting (7.2).
@@
  • If $Y$ is the twisted plane, then by example 6.8 the $2$-dimensional irreducible components of $F$ have degree $1$ and $4$ with respect the Plücker embedding. Again, this contradicts (7.2).
@@
The above shows that we can assume $\dim F=2$. It follows from proposition 6.7 that the Hilbert polynomial of $F$ with respect to the Plücker embedding is
@@
$0\longrightarrow\mathcal{O}_{F}\longrightarrow\gamma_{\ast}\mathcal{O}_{X}\longrightarrow Q\longrightarrow 0$

### v3

@@ 6.3. The Fano variety of singular cubic hypersurfaces
By lemma˜6.11 it remains to treat the case when $X$ is not normal. We then conclude by proposition˜5.3 and examples˜6.6, LABEL:, 6.4, LABEL: and 6.8. ∎
@@ 6.4. A construction
Let $X$ be a smooth projective variety over $k$, $V$ a finite-dimensional $k$-vector space and
@@
$\textup{T}_{x}\gamma\colon\quad\textup{T}_{x}X\;\longrightarrow\hskip-14.0pt\longrightarrow\;\textup{T}_{\gamma(x)}F.$
@@
As a sanity check, note that for $X=\mathbb{P}^{1}$ the hypotheses of proposition˜6.12 are never fulfilled, hence this is not in contradiction to proposition˜6.1.
@@ 7.1. Main result
    the difference morphism $X\times X\to X-X$ has generic degree $\geqslant 6$, and the sum morphism $X\times X\times X\to A$ has an irreducible fiber of dimension $\geqslant 3$.
@@
Combining this result with corollary˜4.7, we obtain the following classification of subvarieties with Tannaka group $E_{6}$:
@@
Let $X\subset A$ be a smooth irreducible subvariety with ample normal bundle and dimension <g/2<g/2. Then the following are equivalent:
@@
The rest of the section is devoted to the proof of theorem˜7.1. By the Lefschetz principle and [JKLM25, cor. 4.4] we may assume $k=\mathbb{C}$. In this case we already know from section˜4.6 that (1) $\Rightarrow$ (2), and (3) $\Rightarrow$ (1) is clear from the example in ˜Example of the introduction. So the only remaining point is (2) $\Rightarrow$ (3), which gives a numerical characterization of Fano surfaces of smooth cubic threefolds.
@@ 7.2. Strategy of the proof
Before coming to the details, let us outline the main steps in the proof of (2) $\Rightarrow$ (3). The idea is to recover the cubic threefold as in the example in ˜Example. By assumption, the subvariety $X\subset A$ is a surface with
@@
Borrow notation from section˜6.4 with the short exact sequence (6.1) being (7.1) and $V=\operatorname{Lie}A$. Since $X$ is nondivisible, it is not invariant by translation under any nonzero abelian variety, thus the Gauss map $\gamma\colon X\to\operatorname{Gr}_{2}(\operatorname{Lie}A)$ is a finite morphism [Deb95, prop. 3.1]. By interpreting $c_{2}(\mathcal{N})$ as the top Segre class of $T_{X}$, the positivity of $c_{2}(\mathcal{N})$ implies that the morphism $\pi\colon\mathbb{P}(T_{X})\to\mathbb{P}(\operatorname{Lie}A)$ is generically finite onto its image. The bootstrap result is the computation of the degree of $\pi$. To do this, we use that the difference morphism
@@
As in section˜6.4 let $Y\subset\mathbb{P}(\operatorname{Lie}A)$ be the image of $\pi$. Then $Y$ has dimension $3$ and degree
@@
In particular, $\deg\pi=6$ and hence $\deg\gamma\in\{1,3\}$. It remains to show that $Y$ is smooth. For this, we will show $F$ is a smooth surface and apply [AK77, th. 4.2]. We begin with:
Proposition 7.5.
The Fano variety $F$ is an integral surface and the finite morphism $\gamma\colon X\to F$ is birational.
To conclude, a comparison of the Hilbert polynomials with respect to the Plücker embedding of $X$ and of the Fano variety $F$ of lines on $Y$ will show:
@@
The rest of this section is devoted to the proof of lemmas˜7.3, LABEL:, 7.4, LABEL:, 7.5, LABEL: and 7.6.
7.3. The projection to the tangent space: Proof of lemma˜7.3
Let us write $d=d_{Z}$ and $\pi=\pi_{Z}$. Let D:=ZZD:=Z-Z be the image of the difference morphism. Consider the blow-up $\tilde{A}$ of $A$ in 0, the blow-up $\tilde{D}$ of $D$ in 0 and the blow-up $\tilde{P}$ of $P=Z\times Z$ in the diagonal. The normal bundle of the diagonal in $Z\times Z$ is by definition the tangent bundle of $Z$. It follows that the exceptional divisor of $\tilde{P}\to P$ is $\mathbb{P}(T_{Z})$. The exceptional divisor of $\tilde{A}\to A$ is $\mathbb{P}(\operatorname{Lie}A)$ and that of $\tilde{D}\to D$ is the tangent cone $C$ of $D$ at 0. Let $\tilde{d}\colon\tilde{P}\to\tilde{D}$ be the morphism induced by the difference map. The situation is summarized in the following commutative diagram:
@@
To conclude the proof of lemma˜7.3, notice first that the top-left square in diagram (7.3) is cartesian. Indeed, this boils down to the fact that the scheme-theoretic fiber of $d$ in 0 is the diagonal $\Delta_{Z}\colon Z\to Z\times Z$. To see this, notice that the difference morphism $A\times A\to A$ is smooth and the scheme-theoretic preimage of 0 is the diagonal $\Delta_{A}\colon A\hookrightarrow A\times A$ with its reduced structure; the claim then follows because the following square is cartesian:
@@
This follows from upper semicontinuity of the rank of $f_{\ast}\mathcal{O}_{S}$ on the locus in $T$ where $f$ is finite. ∎
We apply the preceding lemma with $S=\tilde{P}$, $T=\tilde{D}$, $f=\tilde{d}$ and $T^{\prime}=C$. The hypotheses are fulfilled because $\tilde{d}$ is generically finite, the scheme-theoretic preimage of $C$ is $\mathbb{P}(T_{Z})$, and the restriction of $\tilde{d}$ to $\mathbb{P}(T_{Z})$ is $\pi$ by lemma˜7.7 hence generically finite by hypothesis. Moreover, the morphism $\tilde{d}$ is generically finite of same degree of $d$ because blow-ups are birational maps. ∎
7.4. The cubic hypersurface: Proof of proposition˜7.4
@@
We start by excluding the case $\deg Y=2$. In this case we have $\operatorname{codim}Y=1$ hence $g=5$ and $Y$ is a $3$-dimensional quadric. To rule out the case, let $F$ be the Fano variety of lines in $Y$ and recall by proposition˜6.12 that the image of $\gamma\colon X\to F$ is a $2$-dimensional irreducible component of $F$. The rank $r$ of the quadric $Y$ is necessarily $3$, 44 or $5$ because $Y$ is integral. We proceed now case by case:
  • If $r=3$, then $Y$ is a cone over a smooth plane conic $Q$ and $F$ is irreducible of dimension $\dim Q+2(5-3-1)=3$ by proposition˜6.3. Contradiction.
  • If $r=4$, then $Y$ is a cone over the rational normal scroll $S(1,1)=\mathbb{P}^{1}\times\mathbb{P}^{1}$ in $\mathbb{P}^{3}$ via the Segre embedding [Har92, p. 285]. By proposition˜6.1 the Fano variety of lines in $S(1,1)$ is $\mathbb{P}^{1}\sqcup\mathbb{P}^{1}$, so $F$ has two irreducible components of dimension $\dim\mathbb{P}^{1}+2(5-4)=3$ by proposition˜6.3. Contradiction.
@@
The inequality $\deg Y\geqslant 1+\operatorname{codim}Y$ implies that $Y\subset\mathbb{P}(\operatorname{Lie}A)$ has codimension $1$ or $2$. It remains to rule out the codimension $2$ case. If this is the case, the variety $Y$ is of minimal degree, hence by [EH87, th. 1] it is either a cone over a smooth rational scroll or a cone over the Veronese surface, that is, $\mathbb{P}^{2}$ embedded in $\mathbb{P}^{5}$ by the second Veronese embedding. Now taking a cone with respect to some linear projection preserves the codimension. This rules out at once the case of the Veronese surface since it has codimension $3$ while $Y$ has codimension $2$. Suppose $Y$ is a smooth rational normal scroll $S=S(a)$ or a cone over such a scroll. By proposition˜6.12 the scheme-theoretic image of the finite morphism $\gamma\colon X\to F$ is one of the irreducible components of $F$. A case by case distinction shows that this is a contradiction. More precisely, the tuple $a=(a_{1},\dots,a_{r})$ satisfies $1\leqslant a_{1}\leqslant\cdots\leqslant a_{r}$ and $a_{1}+\cdots+a_{r}=3$. This leaves us with the $3$ possibilities described in detail in example˜6.2:
@@
Suppose $\dim S=1$. In this case the scroll $S$ is the image of the triple Veronese embedding $\mathbb{P}^{1}\hookrightarrow\mathbb{P}^{3}$. As $S$ contains no lines, proposition˜6.3 implies that $F$ is irreducible of dimension $3$. But $X$ dominates a component of $F$. Contradiction.
@@
Suppose $\dim S=2$. In this case the Fano variety is $\mathbb{P}^{1}\sqcup\{\textup{pt}\}$. By proposition˜6.3 the component of $F$ over $\mathbb{P}^{1}$ has dimension $\dim\mathbb{P}^{1}+2(6-5)=3$ hence $X$ cannot dominate it. Instead again by proposition˜6.3 the component of $F$ over the singleton is a projective plane embedded linearly via the Plücker embedding. If $X$ were to dominate it, we would have $\deg\gamma=45$, contradicting (7.2).
@@
Suppose $\dim S=3$. In this case $Y=S$ and $F\simeq(\mathbb{P}^{1}\times\mathbb{P}^{2})\sqcup\mathbb{P}^{2}$. It follows that $X$ has to dominate the component isomorphic to the projective plane. But this component is linearly embedded via the Plücker embedding. As above, if $X$ were to dominate this component, then $\deg\gamma=45$ which contradicts (7.2).
7.5. The Fano surface and the birationality of the Gauss map: Proof of proposition˜7.5
@@
By proposition˜6.9 we have $\dim F\leqslant 3$, with equality if and only if $Y$ is either the twisted plane in example˜6.8 or a cone over a nonnormal cubic surface or a cone over a smooth cubic plane curve. These cases are excluded by looking at the Fano variety:
@@
  • If $Y$ is a cone over an integral cubic curve, then $F$ is irreducible and has dimension $3$; see example˜6.4.
@@
  • If $Y$ is a cone over a nonnormal cubic surface that is not itself a cone over a nonnormal cubic curve, then the $2$-dimensional irreducible components of $F$ are linearly embedded via the Plücker embedding; see example˜6.6. This would imply $\deg\gamma=45$, which contradicts (7.2).
@@
  • If $Y$ is the twisted plane, then by example˜6.8 the $2$-dimensional irreducible components of $F$ have degree $1$ and $4$ with respect the Plücker embedding; this again contradicts (7.2).
The above shows that we can assume $\dim F=2$. To conclude that $F$ is integral, it suffices to show that the Gauss map $\gamma\colon X\to F$ is birational onto its image: Indeed, assuming this, let $\mathcal{L}$ be the line bundle on $F\subset\operatorname{Gr}_{2}(\operatorname{Lie}A)$ which is the restriction of $\mathcal{O}(1)$ on $\mathbb{P}(\operatorname{Alt}^{2}\operatorname{Lie}A)$ via the Plücker embedding. Then $\mathcal{L}$ and $\gamma^{\ast}\mathcal{L}$ have self-intersection
$\mathcal{L}.\mathcal{L}=45=\gamma^{\ast}\mathcal{L}.\gamma^{\ast}\mathcal{L}.$
The first identity holds by proposition˜6.7, the second is $c_{1}(X)^{2}=45$ since $\gamma^{\ast}\mathcal{L}$ is the canonical bundle of $X$ by definition of $\gamma$. Since $\gamma$ is birational onto its image and $\mathcal{L}$ is ample, this forces $F$ to be irreducible. Proposition˜6.12 then implies that $F$ is generically reduced, hence satisfies condition $R_{0}$ because its singular locus has codimension $\geqslant 1$. But proposition˜6.7 implies that $F$ is a locally complete intersection, thus satisfies condition $S_{1}$. Therefore, $F$ is reduced by [Gro65, 5.8.5].
We now show that $\gamma\colon X\to F$ is birational onto its image. To do this, we mimic the relation between incidence divisors and the canonical bundle of the Fano surface of a smooth cubic threefold [CG72, p. 326]. For any $x\in X$, consider the line $L_{x}:=\mathbb{P}(T_{x}X)$ in $\mathbb{P}(\operatorname{Lie}A)$. Inside $X\times X$ consider the locus of couples $(x,y)$ with $L_{x}\cap L_{y}\neq\varnothing$ and let
$D\;\subset\;X\times X$
be the union of its top dimensional components. For $x\in X$, let $D_{x}=p(q^{-1}(x))$ be the fiber of the second projection $q\colon D\to X$, considered as a subvariety of $X$ via the first projection $p\colon D\to X$. By [Ran81, lemma II.12], the nondegeneracy of $X$ implies $D_{x}\neq X$. On the other hand, since $\pi\colon\mathbb{P}(T_{X})\to Y$ is generically finite of degree $6$ and $\deg(\gamma)\in\{1,3\}$, there are at least two distinct lines $L_{x}$, $L_{x^{\prime}}$ through a generic $y\in Y$. Thus, $D_{x}$ is a divisor on $X$ for generic $x\in X$.
Lemma 7.9.
Let $U\subset X$ be the open subset where $q\colon D\to X$ is flat.
  1. (1)
    There exist $x,y,z\in U$ such that the lines $L_{x},L_{y},L_{z}\subset\mathbb{P}(\operatorname{Lie}A)$ are pairwise distinct and span a plane.
  2. (2)
    For any such $x,y,z$, the sum $D_{x}+D_{y}+D_{z}$ is a canonical divisor on $X$.
Proof.
(1) By assumption, there exists an integral fiber $I\subset X^{3}$ of dimension $\geqslant 3$ of the sum morphism $X^{3}\to A$. We first claim that for generic $(x,y,z)\in I$, the three lines $L_{x},L_{y},L_{z}$ are pairwise distinct: Otherwise the fiber $I$ would be contained in the preimage of the big diagonal $\Delta\subset\operatorname{Gr}_{2}(\operatorname{Lie}A)^{3}$ under the morphism
$\gamma^{3}\colon\quad X^{3}\longrightarrow\operatorname{Gr}_{2}(\operatorname{Lie}A)^{3}.$
Now $\Delta$ has three irreducible components which are permuted under the action of the symmetric group $S_{3}$. Since $I$ is irreducible and stable under the action of $S_{3}$, it must lie in the preimage of the intersection the irreducible components of $\Delta$. This intersection is the small diagonal of $\operatorname{Gr}_{2}(\operatorname{Lie}A)^{3}$ and its preimage via $\gamma^{3}$ has dimension $2$ because $\gamma$ is finite and $\dim X=2$, thus contradicting $\dim I\geqslant 3$.
We next claim that for any $(x,y,z)\in I$ the lines $L_{x},L_{y},L_{z}$ are coplanar. Since coplanarity is a closed condition, it suffices to show this for generic $(x,y,z)\in I$. The tangent map of the sum morphism $X^{3}\to A$ has generic rank $\leqslant 3$ on $I$ because it contracts $I$. Since the sum map on $A$ induces the sum on tangent spaces, this implies $\dim(T_{x}X+T_{y}X+T_{z}X)\leqslant 3$ for generic $(x,y,z)\in I$. Passing to projective spaces this means that $L_{x},L_{y},L_{z}\subset\mathbb{P}(\operatorname{Lie}A)$ are coplanar.
Finally, we claim that $I\subset X^{3}$ dominates $X$ via each projection: Since $I$ is irreducible and stable under permutations, we would otherwise have $I=C^{3}$ for a curve $C\subset X$, but then the image of $I$ under the sum map would be $C+C+C\subset A$, which is not a point.
(2) Let $P$ be the plane in $\mathbb{P}(\operatorname{Lie}A)$ spanned by $L_{x}$, $L_{y}$ and $L_{z}$ and $H\subset\operatorname{Gr}_{2}(\operatorname{Lie}A)$ the subvariety of points in $\operatorname{Gr}_{2}(\operatorname{Lie}A)$ corresponding to lines meeting $P$. The subvariety $H$ is a hyperplane section of the Plücker embedding. The nondegeneracy of $X$ implies $\gamma(X)$ meets any such hyperplane properly [Ran81, lemma II.12], hence $\gamma^{-1}(H)$ is a divisor linearly equivalent to $\gamma^{\ast}\mathcal{L}$, which is isomorphic to the canonical bundle of $X$. By construction, we have
$P\cap Y=L_{x}\cup L_{y}\cup L_{z},$
hence $\gamma^{-1}(H)=D_{x}\cup D_{y}\cup D_{z}$, which concludes the proof. ∎
Corollary 7.10.
For $U\subset X$ as above and $x,y\in U$, we have $D_{x}.D_{y}=5$.
Proof.
By design, the divisors $D_{x}$ form a flat family over $U$, thus the intersection number $n=D_{x}.D_{y}$ is independent of $x$ and $y$. For $x,y,z\in U$ such that the lines $L_{x}$, $L_{y}$ and $L_{z}$ are pairwise distinct and span a plane, we thus have
$9n=(D_{x}+D_{y}+D_{z})^{2}=c_{1}(X)^{2}=45,$
because $D_{x}+D_{y}+D_{z}$ is a canonical divisor by lemma˜7.9. Therefore $n=5$. ∎
We can now prove that $\gamma$ is birational onto its image: The divisors $D_{x}$ are the pullback of Weil divisors on $\gamma(X)$ because the incidence relation defining them comes from $\operatorname{Gr}_{2}(\operatorname{Lie}A)$. The projection formula then implies that $\deg(\gamma)$ divides $D_{x}.D_{y}$ for $x,y\in U$. Since $\deg(\gamma)\in\{1,3\}$ and $D_{x}.D_{y}=5$, it follows that $\deg(\gamma)=1$. ∎
7.6. Smoothness: Proof of proposition˜7.6
@@
It follows from proposition˜6.7 that the Hilbert polynomial of $F$ with respect to the Plücker embedding is
@@
$0\longrightarrow\mathcal{O}_{F}\longrightarrow\gamma_{\ast}\mathcal{O}_{X}\longrightarrow Q\longrightarrow 0.$